OpenStudy (anonymous):

Solve for h:

OpenStudy (anonymous):

\[\frac{ 3 }{ 7 }h = 9\]

OpenStudy (anonymous):

ok first u need to get h by its self what do u do then

OpenStudy (anonymous):

or how do u get that?

OpenStudy (anonymous):

its 21

OpenStudy (anonymous):

multiply 7 on both sides?

OpenStudy (anonymous):

no u have to divide 3/7

OpenStudy (anonymous):

ohh okay so \[\frac{ 3 }{ 7 } \div \frac{ 3 }{ 7 }h= 9 \div \frac{ 3 }{ 7 } \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then turn 9 into a fraction

OpenStudy (anonymous):

so then it would be h = \[\frac{ 9 }{ 1 } * \frac{ 7 }{ 3 }\]..?

OpenStudy (anonymous):

okay I got it now thanks for your help :)